Toga Radar Science Log camex – Key West, Florida 17 –8 –01 14: 05 utc



Download 230.93 Kb.
Page1/6
Date conversion29.04.2018
Size230.93 Kb.
  1   2   3   4   5   6
Toga Radar Science Log


CAMEX – Key West, Florida

17 –8 –01 14:05 UTC: Clear day this morning, a few spotty clouds over the ocean. Nothing significant. Just little scattered cells about 100km south, moving west in a straight line. 1 little cell about 20km north or the radar, also moving west in a straight line. NSD
15:46 UTC: A few light showers in the viewable area. Cloud-cover is about 30%. An east-west aligned group of showers, about 40dbz max, is about 100 to 150 km to the south. JRL
16:37 UTC: A large cell has developed right on top of the radar. The cell is rather large and moving to the Southwest. There are also scattered cells about 100 to 150km south of the radar that are moving in the same direction, with reflectivity values in the 18 to 50dBz range. NSD

19:16 UTC: The “Radiate On “ switch has gone out in the Transmitter. I don’t have a replacement for it right now, so I will have to go down to N-Pole to get another one. They have instructed me to shutdown operations for today, and fix the switch tomorrow morning. NSD

21-8-01 15:35 UTC: Finally got the radar working again after 2 days, Now we are up and running. Scattered cells to the northwest moving north at this time. Reflectivity is in the 18 to 60dBz range. At about 100 to 200km away from the radar. Right above us it is overcast, but nothing significant. There was a large cell moving north from the south, but it has dissipated at this time. NSD

16:36 UTC: Area of convection in the northwest quadrant becoming more organized. Two distinct bands are noticeable. The band farthest from the radar (200km) has higher reflectivity values (50 – 60 dBz). Second band (150 km from the radar) is much weaker (30 – 40 dBz). Some isolated cells also located northeast through south. Overcast skies still observed overhead with scattered CU in the low levels. ETA
17:31 UTC: Thunderstorm overhead with light to moderate rain and occasional lightning. Scheduled the KAMP_NEAR task. Areas of convection in the northwest still moving northwest. The entire lower Keys are covered with convection with reflectivity values in the 34 – 44 dBz range. Cells are moving northwest. ETA
18:30 UTC: Thunderstorms previously mentioned moved northwest. Another convective band is overhead moving northwest. Moderate to heavy rain with cloud to ground lightning has been observed. Reflectivity values 34 –44 dBz. Most of the convective region previously located 150 – 200 km from the radar has moved out of radar range or has significantly decreased in intensity. ETA


19:25 UTC: Thunderstorm ended. The convective cells organized into a rather large band extending from the southwest to northeast into extreme southern Florida. The convective band is approximately 300 km in length and 75 km wide. At this time, most reflectivity values fall in the 40 – 50 dBz range. One cell along the 300-310 radial has higher reflectivity values (greater than 60 dBz). This cell was the one from the first wave that moved over the area at 17:31. It increased in intensity as the storms from the second wave (18:30 intersected) intersected with it. ETA

21:15 UTC: The convective band continues to move to the northwest. Another band appears to be trying to organize to the east and southeast. Another region of convection is located in Cuba and is slowly moving westward. Running the SURVEILLANCE and the NORTH_FAR tasks. ETA

22:30 UTC: NO significant changes. ETA
22-8-01 00:45 UTC: There is a band to the east spanning from 75 to 150 degrees breaking up as it moves westward. Land breeze looking cell persistent along the north Cuban coast. Persistent small cell at about 100 km to the northwest traveling south/southeast. Still lots of rain over Florida mainland. CDR + KLM
01:45 UTC: Widespread showers along North Cuban coast. There are small clusters of cells at 125 degrees. Weak cells at 45 degrees. Showers continuing over mainland of Florida. A cell with reflectivity values of 58 dBz popped out of the small previous cell at 325 degrees and 125 km. CDR + KLM
03:33 UTC: Showers along N. Cuba are diminishing. The cell at 325 degrees is now at 150 km, but has broken up into 2 to 3 cells, each with reflectivity values at about 55 dBz. Showers to the east and over the mainland of Florida are also dissipating. CDR
05:00 UTC: Showers along N. Cuba are away from mainland and mainly to the northeast. There is a small cell moving toward us but still 50 km away. The cells at 325 degrees have dissipated and are now 150 to 300 km away. Rain over Florida is insignificant. The evening is fairly calm, but still overcast. CDR

06:00 UTC: Showers over NNE Cuba continue to slowly dissipate, while a few small convective cells have popped up to our east ranging from 075 degrees (200 km) to 100 degrees (100 km). Little cluster of cells keeps creeping up from the SSE, currently about 20km away, no lightning seen yet. S FL is now mostly dry and there appears to be a few convective cells riding some linear body from 100 – 250 km rotating from 315 to 308. Have performed a handful of RHI scans over the last two hours. KLM



07:00 UTC: Light showers that began around 06:40 UTC ended about 5 minutes ago. Showers hanging to the immediate (15 – 75 km) S and SE appear to be mostly stratiform now. No lightning ever observed. Convective cluster at 085 (~180 km) more consolidated. Showers N and NE of Cuba almost gone. A couple of cells continue to pop up along CCW rotating linear body now at 295. Pretty quiet. KLM
08:15 UTC: Little going on now. Showers to our immediate S and SE have dissipated and the only significant returns are progressing our way along a 90 radial, currently about 150km away. KLM
09:10 UTC: Almost all radar returns are between 75 and 140. Previous convective area to the E now producing relatively weak returns beyond 200 km. A couple cells at about 105 showing greater than 34 dBZ, otherwise little to report. KLM
10:20 UTC: A few convective cells still hanging on as we see a little brightening of the sky. To either side of the 130 azimuth, up to five small 30-40 dBz cells remain. Further to their N (at 090) a couple of lingering stratiform showers continue to decay. KLM
11:43 UTC: A small area of relatively weak echoes (34 – 37 dBz) to the east and southeast. Other than that, there is nothing significant to report. TOGA is OPR; currently running the KAMP_FAR and the SURVEILLANCE tasks. ETA

12:57 UTC: No significant changes since the last entry. Set the workstation clock up 10 sec. To agree with the Satellite time, so that we can have accurate comparisons. NSD

14:02 UTC: A few isolated cells are within 50 km of radar site. Cells to the north are moving southwest while those to the south are moving west. Beyond 100 km there are two regions of convection, one to the northwest moving southwest. The other region is in the southeast quadrant. The reflectivity values associated with this region (23 – 37 dBz) are much smaller than those associated with the storms in the northwest (50 – 60 dBz). Still running the SURVEILLANCE and the KAMP_FAR tasks. ETA
15:32 UTC: Scattered cells have appeared in the Northwest over an area of about 50 to 250km away from the radar. There is also a larger cell forming near Miami, about 200 to 300km Northeast of the radar. Reflectivity is in the 18 to 50dBz range. There is also a large cell forming directly South of the radar at about 25 to 50km. The reflectivity in this cell is only in the 12 to 21dBz range at this time, but, I expect it to grow as the day goes by. NSD


16:46 UTC: The cells that were to the south have dissipated. More convection has developed in the Miami area with highest reflectivity values in the 37 – 50 dBz range. The cluster of convection to the northwest still persists although there has been a decrease in coverage. The higher reflectivity values are in the 50 – 55 dBz range. ETA
18:30 UTC: No significant changes since last entry, although there are two little cells forming over Cuba. Reflectivity values are in the 18 to 40 dBz range. NSD

19:20 UTC: There was one isolated rain shower overhead that you can not see on the radar unless you zoom to the 75km view. The shower only lasted for approximately 20 minutes. Most of the convection over the water has dissipated. Southern Florida has a cluster of cells with reflectivity values in the upper 50dBz range. Still running the SURVEILLANCE and the KAMP_FAR tasks. ETA


20:50 UTC: Nothing significant to report. ETA
22:23 UTC: A few cells developed along an old outflow boundary from the storms that were in extreme southern Florida earlier in the day. These new cells are within 30 km of the radar and extend from the east to southwest. Currently running the SURVEILLANCE and the KAMP_FAR tasks. ETA
23:54 UTC: A strong cell is forming over Cuba now, and there are two cells forming to the Southeast, moving Southwest. Reflectivity values are in the 18 to 50dBz range. NSD
23-08-2001 01:01 UTC: Neil changed out data tapes at 00:00z, back in service and scanning (& recording) by 00:20z. As for any wx, the arial coverage of the convection over C and SC Cuba is rapidly diminishing though a small area of ~45dBZ still exists. Elsewhere, several small, unorganized cells exist between 090 and 145 ranging from 35 km to 200 km. A small cell shows up at about 045 (180 km) and a couple more over SW FL mainland at 340 and 250 km. All of these cells have a max reflectivity around 36-42 dBz. KLM + CDR
02:34 UTC: a couple of cells popped up to the E and SE, including on within about 5-10 km. Upper 40dBZ echo at 090 and 105 km, another 40dBZ cell at 125 and a third small cluster at about 145.

Note: While attempting RHI of nearby cell, first RHI at 096.3 rose to an elevation of 92.1 and after sweeping to 140.1 for second RHI scan, remained stuck at 92.1 elev. Mike attempted to bring down antenna manually to see if we could proceed: unsuccessful. Then called Neil.

We shut down Surveillance and KAMP_FAR in IRIS software and Mike climbed up into the dome to manually push the antenna down. Came back and first attempt to restart scanning failed due to EL ERROR indicator still on the RCP02 (box on top of rack #9). Once this was reset, scanning resumed as normal. Performed surveillance scan at 03:03 UTC and next full volume scan began at 03:12 UTC.





  1   2   3   4   5   6


The database is protected by copyright ©hestories.info 2017
send message

    Main page